On its maiden voyage in July 1952, the liner United States won the coveted Blue Ribbon for the fastest crossing of the Atlantic from New York to Cornwall, U.K. The trip took 3 days 10 hours 40 min at an average speed of 34.5 knots (65.5 km/h). This was 10 h 2 min less than the 14-year-old record held by the Queen Mary. What was the average speed of the Queen Mary


1
Expert's answer
2022-01-18T15:13:03-0500

"v_{av}=v\cdot\frac {t_1}{t_1+t_2}=58.4~\frac{km}h."